# Yuehao Bai

### v1

@@ 1 Introduction
The literature on stratified block randomization dates back to at least Fisher (1935). For the case in which $\ell=1$ or $k-\ell=1$, the variance estimator in Imai (2008) has been extended to different settings by Imai et al. (2009), Pashley and Miratrix (2021), and Zhu et al. (2024). Fogarty (2018b) studies regression adjustment for matched pair designs, and Liu and Yang (2020) studies regression adjustment for stratified designs primarily for the case where $\min\{\ell,k-\ell\}>1$. Ding (2017) considers randomization inference for matched pairs, alongside other randomization schemes. We note that all of these papers employ a design-based framework as we do in this paper. As a consequence, their analyses differ from work which studies inference in finely stratified experiments from a superpopulation perspective (Bai, 2022; Bai et al., 2022, 2024a, 2024b, 2024d; Jiang et al., 2024; Cytrynbaum, 2024a, b; Bai et al., 2025); see also Abadie and Imbens (2008), who consider superpopulation estimation of the variance in matched pair designs, conditional on the covariates, in an alternative sampling framework where pairs are sampled instead of the units.
@@ 3.3 A Novel Paired-Strata Estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$
Building on our earlier work on super-population approaches to inference for finely stratified designs (Bai et al., 2022, 2024a, 2024b, 2024d, 2025), we now propose a novel paired-strata estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$, primarily for settings where $\min\{\ell,k-\ell\}=1$. For simplicity, we pair adjacent strata together, so the pairs are given by $\{(\lambda_{2j-1},\lambda_{2j}):1\leq j\leq\lfloor m/2\rfloor\}$, but we emphasize that this pairing should be understood as being obtained after permuting the indices of the strata $\{\lambda_{j}:1\leq j\leq m\}$. In practice, it will be desirable to do this in a way that ensures “similar” strata are adjacent. With this is mind, our estimator is defined as follows:
@@ 4 Comparison with Other Variance Estimators
For some random variable $\tilde{W}:=(\tilde{Y}(1),\tilde{Y}(0),\tilde{X})\sim Q$ with $\max_{d\in\{0,1\}}E_{Q}[\tilde{Y}^{2}(d)]<\infty$, $W^{(n)}$ and $\Lambda_{n}$ satisfy the following requirements:
@@
following similar arguments to those used in Lemma C.2. in Bai et al. (2024d). In words, (13) requires that the average squared distances of the covariate values in a paired stratum converges to zero asymptotically. This property is guaranteed by specific matching algorithms under appropriate assumptions; see, for instance, Bai et al. (2022) and Cytrynbaum (2024b) for details. Assumption 4.2 introduces similar conditions to Assumption 4.1 but also requires that these conditions hold for the covariate vectors $X^{(n)}$. Note in particular that $E_{Q}[\tilde{X}\tilde{Y}(d)]=E_{Q}[\tilde{X}E_{Q}[\tilde{Y}(d)|\tilde{X}]]$ by the law of iterated expectations.
@@
It thus remains to construct a consistent estimator of the final term. For this purpose, it can also be shown under Assumption 4.1 that
@@ 5 Simulations
$E_{Q}[\operatorname{Var}_{Q}[\tilde{Y}(d)|\tilde{X}]]=E_{Q}[\tilde{Y}(d)^{2}]-E_{Q}[E_{Q}[\tilde{Y}(d)|\tilde{X}]^{2}]=\operatorname{Var}_{Q}[\tilde{Y}(d)]+E_{Q}[\tilde{Y}(d)]^{2}-E_{Q}[E_{Q}[\tilde{Y}(d)|\tilde{X}]^{2}]~{}.$

### v2

@@ 1 Introduction
The literature on stratified block randomization dates back to at least Fisher (1935). For the case in which $\ell=1$ or $k-\ell=1$, the variance estimator in Imai (2008) has been extended to different settings by Imai et al. (2009), Pashley and Miratrix (2021), and Zhu et al. (2024). Fogarty (2018b) study regression adjustment for matched pair designs, and Liu and Yang (2020) studies regression adjustment for stratified designs primarily for the case where $\min\{\ell,k-\ell\}>1$. Ding (2017) considers randomization inference for matched pairs, alongside other randomization schemes. We note that all of these papers employ a design-based framework as we do in this paper. As a consequence, their analyses differ from work which studies inference in finely stratified experiments from a superpopulation perspective (Bai, 2022; Bai et al., 2022, 2024a, 2024b, 2024d; Jiang et al., 2024; Cytrynbaum, 2024a, b; Bai et al., 2025); see also Abadie and Imbens (2008), who consider superpopulation estimation of the variance in matched pair designs, conditional on the covariates, in an alternative sampling framework where pairs are sampled instead of the units.
@@ 3.3 A Novel Paired-Strata Estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$
Building on our earlier work on super-population approaches to inference for finely stratified designs (Bai et al., 2022, 2024a, 2024b, 2024d, 2025), we now propose a novel estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$, primarily for settings where $\min\{\ell,k-\ell\}=1$. Our estimator is constructed by pairing together strata; for simplicity, we pair adjacent strata together, so the pairs are given by $\{(\lambda_{2j-1},\lambda_{2j}):1\leq j\leq\lfloor m/2\rfloor\}$, but we emphasize that this pairing should be understood as being obtained after permuting the indices of the strata $\{\lambda_{j}:1\leq j\leq m\}$. In practice, it will be desirable to do this in a way that ensures “similar” strata are adjacent. With this is mind, our estimator is defined as follows:
@@ 4 Comparison with Other Variance Estimators
For some random variable $\tilde{W}:=(\tilde{Y}(1),\tilde{Y}(0),\tilde{X})\sim Q$ with $\max_{d\in\{0,1\}}\allowbreak E_{Q}[\tilde{Y}^{2}(d)]<\infty$, $W^{(n)}$ and $\Lambda_{n}$ satisfy the following requirements:
@@
following similar arguments to those used in Lemma C.2 in Bai et al. (2024d). In words, (13) requires that the average of the squared distances of the covariate values in a paired stratum converges to zero asymptotically. This property is guaranteed by specific matching algorithms under appropriate assumptions; see, for instance, Bai et al. (2022) and Cytrynbaum (2024b) for details. Assumption 4.2 introduces similar conditions to Assumption 4.1 but also requires that these conditions hold for the covariate vectors $X^{(n)}$. Note in particular that $E_{Q}[\tilde{X}\tilde{Y}(d)]=E_{Q}[\tilde{X}E_{Q}[\tilde{Y}(d)|\tilde{X}]]$ by the law of iterated expectations.
@@
It thus remains to construct a consistent estimator of the final term and subtract it from $\hat{\tau}^{2}_{n}$. For this purpose, it can also be shown under Assumption 4.1 that
@@ 5 Simulations
$E_{Q}[\operatorname{Var}_{Q}[\tilde{Y}(d)|\tilde{X}]]$ $=E_{Q}[\tilde{Y}(d)^{2}]-E_{Q}[E_{Q}[\tilde{Y}(d)|\tilde{X}]^{2}]$
$=\operatorname{Var}_{Q}[\tilde{Y}(d)]+E_{Q}[\tilde{Y}(d)]^{2}-E_{Q}[E_{Q}[\tilde{Y}(d)|\tilde{X}]^{2}]~{}.$
@@ 2. Model 2:
6 Recommendations for Empirical Practice
Based on our theoretical results as well as the simulation study above, we conclude with some recommendations for empirical practice. Overall, we argue that our novel estimator $\hat{V}_{n}$ is well suited for performing design-based inference on the average treatment effect, particularly in settings with $\min\{\ell,k-\ell\}=1$. Since the magnitude of the bias of this estimator depends explicitly on the quality of the pairing of the strata, we recommend that practitioners form the strata pairs by applying an optimal non-bipartite matching algorithm (see, for instance, Greevy et al., 2004) to the stratum-level covariate means $\bar{X}_{j}=\frac{1}{k}\sum_{i\in\lambda_{j}}X_{i}$; this algorithm is available in the R package nbpMatching.
An interesting feature of our results relative to our findings in prior work on the analysis of randomized experiments from a super-population perspective (Bai et al., 2022, 2024a, 2024d, 2025) is that, by appealing directly to Theorem 3.2, asymptotically valid inference on $\Delta_{n}$ is possible under minimal assumptions on the nature of the matching procedure. In contrast, inference on the average treatment effect from a super-population perspective seems to require an assumption like (13). However, we note that inferences based on the design-based estimator $\hat{V}_{n}$ are not guaranteed to be valid if we view the sample as being drawn from a larger (finite or super-) population. In such cases, it can be shown that the super-population variance estimators proposed in Bai et al. (2022) (see also Bai et al., 2024d, 2025) remain valid in the limiting thought experiment of Section 4.